# Quantum Knowledge Distillation for Large Language Models

### v1

@@ 4.2.2 Quantum student model
$\left|{\psi_{{\text{ansatz}}}^{1}}\right\rangle=\prod\limits_{i=1}^{n}{{U_{Y}}}({\phi_{i}})\prod\limits_{i,j=i+1}^{n}RZZ({\delta_{i,j}})\prod\limits_{i=1}^{n}{{U_{Z}}}({\lambda_{i}})\prod\limits_{i,j=i+1}^{n}CNOT({q_{i}},{q_{j}})\left|{{\psi_{{\text{enc}}}}}\right\rangle,$ (23)
These operations constitute a single layer of the quantum circuit. Following this, there are $p-1$ additional identical quantum layers, and the output quantum state $\left|\psi_{\text{ansatz}}^{k}\right\rangle$ is measured to obtain the result $m=\text{Measure}(\left|\psi_{\text{ansatz}}^{k}\right\rangle)$. The measurement results form a real-valued vector, and by applying a softmax function, the final probabilities representing the features of the quantum student model are obtained.
4.2.3 Distillation process
After obtaining the outputs from both the LLM teacher model and the quantum student model, QD-LLM need to optimize the quantum student model’s output based on the LLM teacher model and the true labels. To measure the discrepancy between the quantum student model’s output and the LLM teacher model’s output, we employ both KL divergence and JS divergence. KL divergence quantifies the difference between two probability distributions and is commonly used in knowledge distillation tasks [52]. The symmetry of JS divergence ensures a balanced evaluation of the differences between the two models, avoiding bias toward either one, making it a fairer measure of divergence [53].
@@ 4.2.3 Distillation process
$\mathbb{L}(\theta)=\lambda_{1}\mathop{\mathbb{E}}\limits_{x\sim p_{x}}\left[\underbrace{\sum\nolimits_{i}f_{i}(x)\log\frac{f_{i}(x)}{\mathbf{q}_{\theta,i}(x)}}_{\mathbb{L}_{KL}\text{ part}}+\underbrace{\frac{1}{2}\left(\sum\nolimits_{i}f_{i}(x)\log\frac{f_{i}(x)}{m_{i}(x)}+\sum\nolimits_{i}\mathbf{q}_{\theta,i}(x)\log\frac{\mathbf{q}_{\theta,i}(x)}{m_{i}(x)}\right)}_{\mathbb{L}_{JS}\text{ part}}\right]+\lambda_{2}\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}x\sim p_{x}\\ y\sim\tilde{p}(\cdot|x)\end{subarray}}\left[\underbrace{-\sum\nolimits_{i}y_{i}\log\mathbf{q}_{\theta,i}(x)}_{\mathbb{L}_{CE}\text{ part}}\right],$ (24)
@@
4.2.4 Training
During the entire distillation process, the goal of the QD-LLM approach is to minimize the loss function, i.e., argmin ${\text{argmin }}\mathbb{L}$. By minimizing this total loss function, we can adjust the parameters of the quantum student model so that its output becomes increasingly closer to both the LLM teacher model and the true labels. As the value of $\mathbb{L}$ decreases, it indicates that the performance of the quantum student model is approaching that of the LLM teacher model in the specific domain, eventually allowing the quantum student model to complete tasks independently.
4.3 Inference
@@
5 Experiments
In this section, we present a comparative analysis of QD-LLM and existing methods. To minimize the randomness inherent in the optimization process, each solving procedure is repeated five times, and the average of these five values is reported as the result. All experiments were conducted on NVIDIA GeForce RTX 4090 GPUs, and based on the MindSpore Quantum [54] framework.
5.1 Settings
Datasets
To comprehensively evaluate the method, we selected three classic or recent research hotspot tasks in computational linguistics: 1. Emotion analysis [55], 2. Hiding detection [56], and 3. Thematic analysis [57]. Emotion analysis involves understanding and extracting emotional information from text, aiding in the analysis of public opinion and intentions [58]. Hiding detection refers to concealing sensitive information within normal media, making it difficult for unauthorized individuals to detect and extract [59]. Detection of such hiding is critical for preventing misuse of this technology and safeguarding national security [60]. Thematic analysis aims to uncover the implicit themes in text, enabling theme classification and improving the effectiveness of applications such as question-answering systems [57]. The datasets contain an equal number of samples per category and are randomly divided into training, validation, and testing sets in a 6:2:2 ratio, with specific details presented in Table 1.
@@ 5.1 Settings
Baselines
The baselines in this paper are divided into two parts: compression-task baselines and related-task baselines related to the three datasets mentioned above. Compression-task baselines: 1. TinyBERT [61], 2. MiniLLM [19], 3. DistilBERT [14], 4. PKD [62]. Related-task baselines: 5. TextRCNN [63], 6. HiTIN [64], 7. LSFLS [65], and 8. BERT-QTC [45].
  • 1.
    TextRCNN [63] proposes a hierarchical text classification global model that better captures label dependencies and interactions between text and label spaces, achieving significant results.
  • 2.
    HiTIN [64] employs tree isomorphism techniques to enhance hierarchical text classification by accurately mapping label hierarchies, substantially improving classification performance.
  • 3.
    LSFLS [65] presents a low-shot learning framework for text steganalysis, facilitating detection of hidden information with few labeled samples. It enhances adaptability to various steganographic methods, improving performance in low-data scenarios.
  • 4.
    BERT-QTC [45] treats quantum circuits as feature extractors for text classification in heterogeneous computing environments, providing a new perspective on text classification tasks.
Simulation
As we all know, most existing QNNs algorithms in the NISQ era still rely on classical computers to simulate quantum circuits, as current quantum hardware lacks sufficient qubits and fault-tolerant capabilities. Thus, in our experiment, we use classical computer to simulate QD-LLM. Detailed information about the quantum part of the simulation experiment can be found in Table 2.
@@
Hyperparameters
For the teacher models, we selected some open-source LLMs, including BLOOMZ-1.1B, BLOOMZ-3B, OPT-6.7B, LLaMA2-7B, and LLaMA3-8B, with the rank $r$ of LoRA set to 200. In the student model, the learning algorithm used is Adam [66], with an initial learning rate of 0.06, guidance weight $w$ at 0.9, and the number of epochs set to 10, with a batch_size of 8.
5.2 Comparison with compression-task baselines on parameters
@@
TinyBERT [61] 14.35M* 27.49%
MiniLLM [19] 33.36M 63.91%
DistilBERT [14] 52.20M 100.00%
PKD [62] 52.20M 100.00%
@@
5.3 Comparison with compression-task baselines on time cost
@@
Figure 3: Comparison with original LLMs and knowledge distillation baselines on the knowledge distillation time and inference time. The horizontal axis represents the datasets, the vertical axis represents the time cost, and its unit is second.
According to the results in Figure 3, we can see that in the three datasets, the proposed QD-LLM has shorter training and inference time than the existing knowledge distillation baseline method. This shows that the QD-LLM simulated by classical computers has faster processing capabilities than the baseline method, which also proves that quantum computing has immeasurable potential in knowledge distillation. Furthermore, the difference in inference time of tasks comes from the length of the text. Since the length of the text determines the sequence length tensor in the input, teacher network, and quantum student network, it increases exponentially with the increase of length. Therefore, datasets with long texts are more resource-intensive.
5.4 Comparison with compression-task baselines on performance
@@
In addition, to balance the performance and the cost, that is, the required model parameters or knowledge distillation time, we also give two comprehensive indicators: 5. Acc/Parameters and 6. Acc/Tkd (knowledge distillation time). In other words, the advantages of the method cannot be judged only by the speed of knowledge distillation or performance, but also by comprehensive consideration of these aspects. These specific performances are shown in Table 4, Table 5, and Table 6.
Table 4: Comparison on performance in Emotion analysis dataset. $\uparrow$” means the higher the corresponding value, the better. “Bold” represents the best result. “ *” represents the suboptimal result. “Acc/Param” represents the benefits that each parameter can bring to the Acc. “Acc/Tkd” represents the benefits that one second distillation time can bring to the Acc. “e-$\eta$” represents $1\times 10^{\eta}$ in scientific notation.
@@
Table 5: Comparison on performance in Hiding detection dataset. $\uparrow$” means the higher the corresponding value, the better. “Bold” represents the best result. “ *” represents the suboptimal result. “Acc/Param” represents the benefits that each parameter can bring to the Acc. “Acc/Tkd” represents the benefits that one second distillation time can bring to the Acc. “e-$\eta$” represents $1\times 10^{\eta}$ in scientific notation.
@@
Table 6: Comparison on performance in Thematic analysis dataset. $\uparrow$” means the higher the corresponding value, the better. “Bold” represents the best result. “ *” represents the suboptimal result. “Acc/Param” represents the benefits that each parameter can bring to the Acc. “Acc/Tkd” represents the benefits that one second distillation time can bring to the Acc. “e-$\eta$” represents $1\times 10^{\eta}$ in scientific notation.
@@
According to the results in Table 4 to Table 6, we can see that in the three datasets, the proposed QD-LLM method has generally excellent performance. In addition, based on the comparison of the two constructed comprehensive evaluation indicators, it can be found that the QD-LLM method can bring excellent performance at a lower cost, which also shows the potential advantages of quantum computing in the field of knowledge distillation. Furthermore, the QD-LLM method performs best on the thematic analysis dataset and second on the hiding detection dataset. Compared with the binary classification, thematic analysis is a more challenging four-class classification problem, which shows that QD-LLM has greater potential in more complex classification tasks.
In conclusion, QD-LLM not only surpasses baselines in memory usage and inference speed but also exhibits outstanding adaptability and efficiency in complex multitasking scenarios. This demonstrates the advantages of our proposed student model in knowledge distillation.
5.5 Comparison with related-task baselines
When compared with leading specific designed baselines across various fields, student model obtained by QD-LLM demonstrates exceptional performance, particularly in the core tasks of sentiment analysis, steganalysis, and topic analysis. Despite the strong performance of baseline methods within their respective domains, the student model obtained by QD-LLM shows the ability to surpass these algorithms, achieving significant improvements in accuracy and F1 scores. This indicates that the QD-LLM-optimized student model not only possesses the capability to independently complete complex tasks but also exhibits excellent cross-task generalization, consistently outperforming existing leading algorithms.
@@ 5.5 Comparison with related-task baselines
The results show that QD-LLM is an effective knowledge distillation method by two aspects:
First, as a knowledge distillation method, QD-LLM showcases outstanding performance. Specifically, it not only achieves significant results in model compression but also excels in memory usage, inference speed, and training efficiency. Compared to other knowledge distillation methods, QD-LLM maintains performance comparable to or even exceeding that of larger models while significantly reducing resource consumption, which is its core advantage.
Second, the student model obtained by QD-LLM is capable of completing tasks independently and exhibits strong generalization across different tasks. Whether in sentiment analysis, steganalysis, or topic analysis, the student model outperforms existing baseline methods, showing notable improvements in accuracy and F1 scores.
5.6 Ablation experiments
5.6.1 Ablation of the different loss functions
We also explored the impact of different loss functions on the performance. We conducted ablation experiments on the loss function of the proposed QD-LLM method in the five LLMs and three tasks in Tables 4 to 6.
In order to clearly and intuitively show the performance benefits brought by different loss functions, we integrated the data and obtained the final results, as shown in Table 7.
Table 7: Ablation of the different loss functions. “CE” is the sum of labels for interactive optimization model, “KL” and “JS” are the sum of teacher models for interactive optimization model, “KL+JS+CE” is the loss function of the complete QD-LLM method. The specific content is shown in “Section 4.2.3”. Here we present the average results in different datasets and LLMs. “Bold” represents the best result.
@@ 5.6.1 Ablation of the different loss functions
KL+JS+CE (QD-LLM) 0.8242 0.8254
The results in Table 7 show that QD-LLM requires different loss functions to comprehensively optimize the quantum student model, so that the quantum student model can be closer to the performance of the LLM teacher model.
5.6.2 Ablation of the different LLMs
In addition, since the distilled LLMs are also a variable in the proposed method, we also conducted ablation experiments under different LLMs. The specific results are shown in Table 8.
@@ 5.6.2 Ablation of the different LLMs
From the results in Table 8, it can be found that the overall performance of the proposed QD-LLM method shows an increasing trend as the scale of the LLM teacher model increases. This is because larger-scale LLMs can provide better knowledge for the quantum student model, so that after fine-tuning, it can more effectively perform reasoning tasks in this field alone. Furthermore, this increasing trend does not always exist, because the scale of the quantum student model is fixed and it has a performance upper limit. When the quantum student model reaches the performance upper limit, it is difficult to improve the performance of the quantum student model only by increasing the scale of LLMs. Therefore, it is necessary to increase the scale of quantum bits and quantum student models in order to further improve their performance.
6 Conclusion
As the scale of LLMs continues to expand, there remains room for optimization in model compression. This paper presents a novel approach to distilling LLMs using quantum algorithms, called QD-LLM. QD-LLM consists of an LLM teacher model and a quantum student model, with the quantum student simulating the distribution of outputs from LLMs and the true results. During the inference phase, the trained quantum student model can efficiently handle various tasks. Experimental results confirm the feasibility of this method, with the quantum student network demonstrating reduced parameter count, lower memory usage, and improved inference efficiency compared to baselines, while also achieving superior model performance. Compared to mainstream knowledge distillation baselines and the original LLMs, QD-LLM significantly reduces the number of training parameters, memory consumption, and inference time while maintaining performance, thereby alleviating the challenges of applying LLMs in resource-constrained environments.
It is noteworthy that QD-LLM has currently only been tested on multiple classification tasks. Due to the limitations of traditional computers in simulating large-scale qubits, it is difficult to demonstrate similar advantages in generative tasks and others. Furthermore, the performance of QD-LLM is not the upper limit of quantum algorithms for compressing LLMs. By increasing the number of qubits, optimizing quantum circuit design, and improving the distillation process, further enhancements in distillation effectiveness may be achieved. In the future, improving the compression effects of LLMs and enhancing the performance of the quantum student network are worthwhile research directions. Furthermore, we hope that QD-LLM can provide valuable insights into exploring the construction of quantum distillation methods for other tasks.
7 Discussion
Although the proposed QD-LLM method can distill large-scale language models well and perform well on multiple datasets, they are still essentially classification tasks. We tried more complex task types, including language understanding and generation tasks, but found that the QD-LLM method performed poorly on these tasks. We believe that this may be due to the characteristics of the generation task requiring a mechanism similar to the LLMs, that is, building a quantum vocabulary to generate quantum states by encoding tokens. However, since existing classical computers have difficulty simulating large-scale quantum bits, it is impossible to effectively encode such a quantum vocabulary.
8 Future Work
The performance of QD-LLM is not unlimited. In other words, increasing the size of the teacher model does not necessarily distill a better QD-LLM. This is because the performance of QD-LLM depends on the nonlinear fitting ability of the quantum student network, which is limited by the number of quantum bits. If the number of quantum bits is not increased and the compression level is only improved by increasing the size of the teacher model, the performance improvement of the quantum student model will become extremely limited. In the next research, we will try to improve the performance ceiling of QD-LLM by changing the structure of quantum neural networks or exploring how to run large-scale quantum bits. In addition, how to achieve efficient encoding under the condition of a limited number of quantum bits so that the quantum network can be competent for generation tasks is the focus of future research.
Acknowledgment

### v2

@@ Results
$\left|{\psi_{{\text{ansatz}}}^{1}}\right\rangle=\prod\limits_{i=1}^{n}{{U_{Y}}}({\phi_{i}})\prod\limits_{i,j=i+1}^{n}RZZ({\delta_{i,j}})\prod\limits_{i=1}^{n}{{U_{Z}}}({\lambda_{i}})\prod\limits_{i,j=i+1}^{n}CNOT({q_{i}},{q_{j}})\left|{{\psi_{{\text{enc}}}}}\right\rangle.$ (13)
These operations constitute a single layer of the quantum circuit. Following this, there are $p-1$ additional identical quantum layers, and the output quantum state $\left|\psi_{\text{ansatz}}^{k}\right\rangle$ is measured to obtain the result $m=\text{Measure}(\left|\psi_{\text{ansatz}}^{k}\right\rangle)$. The measurement results form a real-valued vector, and by applying a softmax function, the final probabilities representing the features of the quantum student model are obtained. We note that the QD-LLM framework is not limited to the specific student circuit used in this study. Variants with deeper layers, alternative encoding strategies, or hardware-efficient ansätze may further enhance performance and generality. We note that the QD-LLM framework is not limited to the specific student circuit used in this study. Variants with deeper layers, alternative encoding strategies, or hardware-efficient ansätze may further enhance performance and generality.
Distillation process
After obtaining the outputs from both the LLM teacher model and the quantum student model, QD-LLM needs to optimize the quantum student model’s output based on the LLM teacher model and the true labels. To measure the discrepancy between the quantum student model’s output and the LLM teacher model’s output, we employ both KL divergence and JS divergence. KL divergence quantifies the difference between two probability distributions and is commonly used in knowledge distillation tasks34. The symmetry of JS divergence ensures a balanced evaluation of the differences between the two models, avoiding bias toward either one, making it a fairer measure of divergence35.
@@
$\mathbb{L}(\theta)=\,$ $\lambda_{1}\mathop{\mathbb{E}}\limits_{x\sim p_{x}}\Bigg{[}\underbrace{\sum\nolimits_{i}f_{i}(x)\log\frac{f_{i}(x)}{\mathbf{q}_{\theta,i}(x)}}_{\mathbb{L}_{KL}}+\underbrace{\frac{1}{2}\left(\sum\nolimits_{i}f_{i}(x)\log\frac{f_{i}(x)}{m_{i}(x)}+\sum\nolimits_{i}\mathbf{q}_{\theta,i}(x)\log\frac{\mathbf{q}_{\theta,i}(x)}{m_{i}(x)}\right)}_{\mathbb{L}_{JS}}\Bigg{]}$
$+\lambda_{2}\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}x\sim p_{x}\\ y\sim\tilde{p}(\cdot|x)\end{subarray}}\left[\underbrace{-\sum\nolimits_{i}y_{i}\log\mathbf{q}_{\theta,i}(x)}_{\mathbb{L}_{CE}}\right],$ (14)
@@
Training and inference
During the entire distillation process, the goal of the QD-LLM approach is to minimize the loss function, i.e., ${\text{argmin }}\mathbb{L}$. By minimizing this total loss function, we can adjust the parameters of the quantum student model so that its output becomes increasingly closer to both the LLM teacher model and the true labels. As the value of $\mathbb{L}$ decreases, it indicates that the performance of the quantum student model is approaching that of the LLM teacher model in the specific domain, eventually allowing the quantum student model to complete tasks independently.
@@
Experiments
In this paper, we present a comparative analysis of QD-LLM and existing methods. To minimize the randomness inherent in the optimization process, each solving procedure is repeated five times, and the average of these five values is reported as the result. All experiments were conducted on NVIDIA GeForce RTX 4090 GPUs . In addition to numerical simulation, we further verify the practicality of QD-LLM through hardware-level implementation. Specifically, we deploy the trained quantum circuits on the Baihua superconducting quantum processor, a 136-qubit device available through the Quafu cloud platform. The experimental procedure and results are detailed below. The code of this article has been open-sourced at https://github.com/Lilingxiao-bupt/QD-LLM.
Simulation Experiments
Although QD-LLM is built on quantum circuit structures, all simulations in this section are performed entirely on classical hardware. Therefore, we treat our method as a quantum-inspired classical algorithm in the simulation experiments. Under this interpretation, QD-LLM is treated purely as a classical algorithm, and its performance can be fairly compared with other classical distillation baselines. All competing models share the same input representation and are constrained to similar model size and training mechanisms. Therefore, comparisons in terms of number of parameters, memory usage, training time, and inference latency are all performed under consistent conditions, ensuring the validity and significance of the observed results.
Settings
Datasets. To comprehensively evaluate the method, we selected three classic or recent research hotspot tasks in computational linguistics that have practical applications: 1. Emotion analysis 27, 2. Hiding detection 28, and 3. Thematic analysis 29. Emotion analysis involves understanding and extracting emotional information from text, aiding in the analysis of public opinion and intentions 36. Hiding detection refers to concealing sensitive information within normal media, making it difficult for unauthorized individuals to detect and extract 37. Detection of such hiding is critical for preventing misuse of this technology and safeguarding national security 38. Thematic analysis aims to uncover the implicit themes in text, enabling theme classification and improving the effectiveness of applications such as question-answering systems 39. The datasets contain an equal number of samples per category and are randomly divided into training, validation, and testing sets in a 6:2:2 ratio, with specific details presented in Table 1.
@@
Baselines. The baselines in this paper are divided into two parts: compression-task baselines and related-task baselines related to the three datasets mentioned above. Compression-task baselines: 1. TinyBERT 40, 2. MiniLLM 12, 3. DistilBERT 41, 4. PKD 42. Related-task baselines: 5. TextRCNN 43, 6. HiTIN 44, 7. LSFLS 45, and 8. BERT-QTC 46.
  1. 1.
    TextRCNN proposes a hierarchical text classification global model that better captures label dependencies and interactions between text and label spaces, achieving significant results.
  2. 2.
    HiTIN employs tree isomorphism techniques to enhance hierarchical text classification by accurately mapping label hierarchies, substantially improving classification performance.
  3. 3.
    LSFLS presents a low-shot learning framework for text steganalysis, facilitating the detection of hidden information with few labeled samples. It enhances adaptability to various steganographic methods, improving performance in low-data scenarios.
  4. 4.
    BERT-QTC treats quantum circuits as feature extractors for text classification in heterogeneous computing environments, providing a new perspective on text classification tasks.
Simulation: As we all know, most existing QNN algorithms in the NISQ era still rely on classical computers to simulate quantum circuits, as current quantum hardware lacks sufficient qubits and fault-tolerant capabilities. Thus, in our experiment, we use the classical computer to simulate QD-LLM. Detailed information about the quantum part of the simulation experiment can be found in Table 2.
@@
Hyperparameters. For the teacher models, we selected some open-source LLMs, including BLOOMZ-1.1B, BLOOMZ-3B, OPT-6.7B, LLaMA2-7B, and LLaMA3-8B, with the rank $r$ of LoRA set to 64. In the student model, the learning algorithm used is Adam 47, with an initial learning rate of 0.06, guidance weight $w$ at 0.9, and the number of epochs set to 10, with a batch_size of 8.
Comparison with compression-task baselines on parameters
@@
TinyBERT 40 14.35M* 27.49%
MiniLLM 12 33.36M 63.91%
DistilBERT 41 52.20M 100.00%
PKD 42 52.20M 100.00%
@@
Comparison with compression-task baselines on time cost
@@
Figure 3: Comparison with original LLMs and knowledge distillation baselines on the knowledge distillation time and inference time. All models are executed on classical hardware, and all reported time and memory values reflect classical computational cost. QD-LLM is simulated as a quantum-inspired classical algorithm, making the comparison with classical baselines fully fair and meaningful. The horizontal axis represents the datasets, the vertical axis represents the time cost, and its unit is seconds.
According to the results in Figure 3, we can see that in the three datasets, the proposed QD-LLM has shorter training and inference time than the existing knowledge distillation baseline method. Although QD-LLM is built upon quantum circuit structures, the simulations are performed entirely on classical hardware. The parameters of the model are classical, the memory consumption is measured in classical bits, and the runtime reflects actual CPU/GPU execution. Therefore, the observed efficiency gains—reduced parameter count, lower memory usage, and faster computation—are genuine and independent of any quantum-specific acceleration. These results suggest that even when treated purely as a classical algorithm, the quantum-inspired design of QD-LLM leads to tangible improvements in computational efficiency. This supports our interpretation of QD-LLM as a quantum-inspired classical algorithm and highlights the potential of quantum design principles to inform classical model compression. Furthermore, the difference in inference time of tasks comes from the length of the text. Since the length of the text determines the sequence length tensor in the input, teacher network, and quantum student network, it increases exponentially with the increase of length. Therefore, datasets with long texts are more resource-intensive.
Comparison with compression-task baselines on performance
@@
In addition to balance the performance and the cost, that is, the required model parameters or knowledge distillation time, we also give two comprehensive indicators: 5. Acc/Parameters and 6. Acc/Tkd (knowledge distillation time). In other words, the advantages of the method cannot be judged only by the speed of knowledge distillation or performance, but also by a comprehensive consideration of these aspects. These specific performances are shown in Supplementary Materials Table 4, Table 5, and Table 6.
[…]
@@
[…]
@@
[…]
@@
According to the results in Supplementary Materials Table 4 to Table 6, we can see that in the three datasets, the proposed QD-LLM method has generally excellent performance. In addition, based on the comparison of the two constructed comprehensive evaluation indicators, it can be found that the QD-LLM method can bring excellent performance at a lower cost, which also shows the potential advantages of quantum computing in the field of knowledge distillation. Furthermore, the QD-LLM method performs best on the thematic analysis dataset and second on the hiding detection dataset. Compared with the binary classification, thematic analysis is a more challenging four-class classification problem, which shows that QD-LLM has greater potential in more complex classification tasks.
In conclusion, QD-LLM not only surpasses baselines in memory usage and inference speed but also exhibits outstanding adaptability and efficiency in complex multitasking scenarios. These results highlight the effectiveness of our quantum-inspired student model and underscore the potential of structured circuit-based distillation as a general strategy for compressing LLMs.
Comparison with related-task baselines
When compared with leading specifical designed baselines across various fields, the student model obtained by QD-LLM demonstrates exceptional performance, particularly in the core tasks of sentiment analysis, steganalysis, and topic analysis. Despite the strong performance of baseline methods within their respective domains, the student model obtained by QD-LLM shows the ability to surpass these algorithms, achieving significant improvements in accuracy and F1 scores, which are shown in Figure 4. This indicates that the QD-LLM-optimized student model not only possesses the capability to independently complete complex tasks but also exhibits excellent cross-task generalization, consistently outperforming existing leading algorithms.
@@
The results show that QD-LLM is an effective knowledge distillation method in two aspects:
  1. 1.
    As a knowledge distillation method, QD-LLM showcases outstanding performance. Specifically, it not only achieves significant results in model compression but also excels in memory usage, inference speed, and training efficiency. Compared to other knowledge distillation methods, QD-LLM maintains performance comparable to or even exceeding that of larger models while significantly reducing resource consumption, which is its core advantage.
  2. 2.
    The student model obtained by QD-LLM is capable of completing tasks independently and exhibits strong generalization across different tasks. Whether in sentiment analysis, steganalysis, or topic analysis, the student model outperforms existing baseline methods, showing notable improvements in accuracy and F1 scores.
Ablation experiments
Ablation of the different loss functions. We also explored the impact of different loss functions on the performance. We conducted ablation experiments on the loss function of the proposed QD-LLM method in the five LLMs and three tasks.
Table 7: Ablation of the different loss functions. “CE” is the sum of labels for the interactive optimization model, “KL” and “JS” are the sum of teacher models for the interactive optimization model, and “KL+JS+CE” is the loss function of the full QD-LLM method. Here we present the average results in different datasets and LLMs. “Bold” represents the best result.
@@
KL+JS+CE (Full QD-LLM) 0.8242 0.8254
@@
To clearly and intuitively show the performance benefits brought by different loss functions, we integrated the data and obtained the final results, as shown in Table 7. The results in Table 7 show that QD-LLM requires different loss functions to comprehensively optimize the quantum student model, so that the quantum student model can be closer to the performance of the LLM teacher model.
Ablation of the different LLMs. In addition, since the distilled LLMs are also a variable in the proposed method, we also conducted ablation experiments under different LLMs. The specific results are shown in Table 8.
@@
From the results in Table 8, it can be found that the overall performance of the proposed QD-LLM method shows an increasing trend as the scale of the LLM teacher model increases. This is because larger-scale LLMs can provide better knowledge for the quantum student model, so that after fine-tuning, it can more effectively perform reasoning tasks in this field alone. Furthermore, this increasing trend does not always exist, because the scale of the quantum student model is fixed, and it has a performance upper limit. When the quantum student model reaches the performance upper limit, it is difficult to improve the performance of the quantum student model only by increasing the scale of LLMs. Therefore, it is necessary to increase the scale of quantum bits and quantum student models in order to further improve their performance.
Real-Device Experiments
To evaluate the real-device feasibility of QD-LLM, we implemented the trained quantum student model on the Baihua superconducting quantum processor via the Quafu cloud platform. Baihua is a 156-qubit chip developed by the Beijing Academy of Quantum Information Sciences, featuring a planar transmon architecture with tunable couplers and OpenQASM-based circuit execution 48.
Figure 5: Real-device performance of QD-LLM. Real-device classification accuracy of the quantum student model, distilled via QD-LLM from three large language models (BLOOMZ-3B, OPT-6.7B, and LLaMA2-7B), across three downstream tasks. Each colored line corresponds to one dataset, with accuracy evaluated on the Baihua superconducting quantum processor.
The quantum circuits used in the real-device experiments were derived from the trained QD-LLM quantum student model obtained during the knowledge distillation stage. These circuits, with fixed parameters, were compiled into OpenQASM format and submitted to the Quafu backend via its API. For each input sample, a variational quantum circuit embedding the feature vector was executed with 10,240 shots. The resulting bitstring measurements were post-processed to compute Pauli-Z expectation values from the 11 selected qubits, forming the quantum embedding that was subsequently passed into a classical classifier. Considering the high computational cost of training on real quantum computers, we selected a subset of representative samples from each classification task for hardware evaluation. This allows us to validate the real-device behavior of QD-LLM circuits while maintaining reasonable execution time and fidelity stability. As shown in Figure 5, the evaluation assesses the classification accuracy of QD-LLM student models distilled from three LLMs on three representative datasets. Despite modest degradation due to hardware noise and finite sampling, the quantum student models exhibit consistent and competitive performance across tasks. Notably, the performance is maintained above 67.85% on both 2-classification and 4-classification tasks, with the highest reaching 73.75%. The overall performance trend remains roughly consistent with the trend presented in the simulation experiments. These findings underscore the robustness and generality of the QD-LLM framework and establish one of the first end-to-end demonstrations of language model distillation circuits deployed on real NISQ hardware.
Conclusion
As the scale and complexity of large language models (LLMs) continue to increase, developing efficient compression strategies remains a critical challenge. In this work, we introduce QD-LLM, a quantum knowledge distillation framework that transfers knowledge from LLMs into variational quantum circuits using the mind quantum 49. During inference, the trained quantum student operates independently on multiple classification benchmarks. In simulation, QD-LLM demonstrates strong performance compared to several classical distillation baselines in the classification task, despite operating entirely on classical hardware. Hence, we can consider the simulation of QD-LLM as a new quantum-inspired classical algorithm, which is a direct simulation of quantum algorithms. Even at the scale of only 11 qubits, QD-LLM still outperforms the classical baseline in multiple dimensions, demonstrating that quantum structural design can guide efficient learning under limited resources. Such emerging simulation results imply that, when the quality of qubits in real hardware is improved, our quantum student model may emerge as a promising candidate for achieving quantum supremacy on practically meaningful tasks, even though the inference accuracy of current hardware is limited. We further deploy the student model obtained from QD-LLM on the Baihua superconducting quantum processor via the Quafu cloud platform. Despite hardware noise and decoherence inherent to the NISQ regime, the quantum student exhibits stable performance, confirming that compressed language models can be effectively executed on contemporary quantum hardware.
It is worth noting that the current implementation of QD-LLM is limited to classification tasks. Due to the difficulty of simulating large-scale quantum systems on classical hardware, it remains challenging to extend this framework to generative tasks, which require sequential token modeling and quantum vocabulary encoding. Moreover, we observe that increasing the size of the teacher model does not guarantee improved student performance, as the capacity of the quantum student is constrained by the number of available qubits and circuit expressivity. Future research should focus on extending QD-LLM to more complex NLP tasks by designing deeper and more expressive quantum circuits, enabling scalable encoding mechanisms, and exploring hybrid schemes that combine quantum and classical modules. Additionally, investigating the performance limits of quantum distillation under increasing quantum resources may yield further insights into the capabilities of quantum-enhanced model compression.
Acknowledgements
We acknowledge the support of Quafu Quantum Cloud Computing Platform from Beijing Academy of Quantum Information Sciences (https://quafu-sqc.baqis.ac.cn/). We also thank Professor Zhangqi Yin and his team at Beijing Institute of Technology for their valuable assistance during the execution of the real-device experiments.
Funding